# Modulate and Reconstruct: Learning Hyperspectral Imaging from Misaligned Smartphone Views

### v3

@@ Abstract
Hyperspectral reconstruction (HSR) from RGB images is a fundamentally ill-posed problem due to severe spectral information loss. Existing approaches typically rely on a single RGB image, limiting reconstruction accuracy. In this work, we propose a novel multi-image-to-hyperspectral reconstruction (MI-HSR) framework that leverages a triple-camera smartphone system, where two lenses are equipped with carefully selected spectral filters. Our configuration, grounded in theoretical and empirical analysis, enables richer and more diverse spectral observations than conventional single-camera setups. To support this new paradigm, we introduce Doomer, the first dataset for MI-HSR, comprising aligned images from three smartphone cameras and a hyperspectral reference camera across diverse scenes. We show that the proposed HSR model achieves consistent improvements over existing methods on the newly proposed benchmark. In a nutshell, our setup allows 30% towards more accurately estimated spectra compared to an ordinary RGB camera. Our findings suggest that multi-view spectral filtering with commodity hardware can unlock more accurate and practical hyperspectral imaging solutions.
@@ 1 Introduction
Hyperspectral imaging provide dense spectral measurements at each spatial pixel forming a 3D cube $\mathbf{I}_{\mathrm{HS}}\in\mathbb{R}^{h\times w\times n}$, where $n\gg 3$. This enables fine-grained analysis of material properties in applications ranging from remote sensing [1], to medical diagnostics [2], to historical preservation [3], to ISP improvement [4, 5], to food quality assessment [6, 7]. However, acquiring such high-dimensional data typically requires hardware expensive, bulky, and often reliant on time-consuming scanning thus fundamentally limiting the usability of hyperspectral imaging in dynamic or consumer settings.
@@
An increasingly studied alternative is hyperspectral reconstruction (HSR): recovering $\mathbf{I}_{\mathrm{HS}}$ from an RGB observation $\mathbf{I}_{\mathrm{RGB}}\in\mathbb{R}^{h\times w\times 3}$ captured via a sensor with spectral sensitivity matrix $\mathbf{S}\in\mathbb{R}^{3\times n}$ under noise $\mathbf{N}$:
$\mathbf{I}_{\mathrm{RGB}}(x,y)=\mathbf{S}\cdot\mathbf{I}_{\mathrm{HS}}(x,y)+\mathbf{N}(x,y).$
@@
This inverse problem is highly ill-posed and advances in deep learning-based HSR [8, 9], reconstructing $\mathbf{I}_{\mathrm{HS}}$ from a single RGB view remains fundamentally limited by the low spectral observability.
@@
Several efforts have aimed to improve this observability. Learned or optimized multispectral filter arrays (MSFAs) [10], end-to-end spectral sensitivity learning [11], and joint sensor-network co-design [12] have all been proposed. However, these methods assume control over sensor hardware and manufacturing thus making them impractical for scalable or consumer-level deployment.
@@
One underexplored but promising path is to leverage the multi-camera systems found in modern smartphones. These devices already include multiple rear cameras with different lenses and spectral sensitivities. In principle, such a setup can be treated as a low-cost, multi-spectral capture system thus capable of observing a scene through multiple, distinct $\mathbf{S}_{i}$ matrices. However, prior works on leveraging multiple cameras [13, 14] do not tackle with unavoidable image misalignment.
@@
Can we turn a multi-camera smartphone into a compact, spectrally diverse imaging system, without altering its internal hardware? Our key insight is that by modulating the spectral response of the auxiliary cameras using carefully chosen external filters, we can create a 9-channel imaging system from off-the-shelf devices. Unlike synthetic MSFAs or custom hardware, our approach requires no internal modification and enables real-time, parallel acquisition. Critically, this setup is easy to deploy, manufacturable at scale, and compatible with existing mobile infrastructure.
@@
We select filters using spectral information loss minimization with respect to a representative hyperspectral distribution. The resulting setup produces spatially misaligned but spectrally rich multi-view data thus posing a new fusion problem that we address via alignment-aware learning. An overview of our physical configuration is shown in Fig. 1, which illustrates how external filters are applied to the auxiliary lenses of a standard smartphone to create a spectrally diverse input set. The combination of low-cost physical augmentation and learning-based reconstruction represents a practical path toward deployable hyperspectral imaging in unconstrained environments. We propose a complete, low-cost pipeline for multi-image-to-hyperspectral reconstruction (MI-HSR) using a filter-modified smartphone. Specifically, the contributions are three-fold:
  • A novel smartphone-based acquisition system that uses two custom spectral filters over auxiliary cameras, converting a consumer-grade smartphone into a 9-channel imaging device. We analyze and justify our filter choices via information-theoretic criteria. To the best of our knowledge, such a configuration has not been previously explored in HSR literature. Our system significantly outperforms RGB-only and naive multi-view baselines.
  • The Doomer dataset, the first benchmark for MI-HSR. It contains 4 captures per scene: three from each of smartphone’s camera and the fourth from the hyperspectral camera.
  • A principled reformulation of transformer-based HSR architecture for our setting, showing that spatial-first attention enables implicit alignment and effective fusion of misaligned inputs across camera viewpoints.
@@ 2 Related Work
Low-cost multispectral imaging. Numerous approaches have aimed to capture multispectral or hyperspectral information without expensive hardware. Early work by Helling et al. [15] employed a grayscale camera with a rotating filter wheel, while Valero et al. [16] used an RGB camera with three interchangeable filters. Oh et al. [14] captured scenes with three different cameras, leveraging the variation in their spectral sensitivities. More recently, Sharma et al. [13] demonstrated that consumer mobile devices with both RGB and NIR sensors can achieve extended spectral capture (400–1000 nm). Although these systems reduce costs, they either involve long capture times or assume no misalignment between successive captures.
Hyperspectral reconstruction from RGB. Traditional HSR methods model spectra using sparse coding [17], dictionary learning [18], or manifold embeddings [19], based on the low-dimensional structure of hyperspectral data and the rarity of metamers [20]. These approaches are computationally efficient but often lack the capacity to incorporate global context from the input image, making them less robust to complex natural scenes. Recent methods based on deep learning have achieved significant advances, particularly those developed through the NTIRE spectral reconstruction challenges [21, 22, 23]. Early deep models used CNNs [9, 24, 25, 26, 27, 28], while newer transformer-based approaches like MST++ [8] and MSFN [29] introduced attention along spectral or spatial dimensions. However, nearly all research relies on synthetic RGB inputs rendered from hyperspectal images (HSIs) using color matching functions (CMFs), assuming perfect alignment and access to camera parameters. These assumptions do not hold in practice, limiting model generalizability. We address these limitations training and evaluating on real-world data with acquisition artifacts and misalignment.
Handling misalignment. Misalignment between inputs and ground truth is a well-known challenge in video and reference-based super-resolution tasks. Optical flow (OF) [30, 31], deformable convolutions [32, 33], and attention mechanisms [34] have been proposed to mitigate this. In our context, misalignment also arises due to ground truth HSIs being not spatially aligned with the RGB input. Zhang et al. [35] addressed this by warping ground truth toward the input using OF, enabling pixel-level evaluation. Elezabi et al. [36] proposed contextual losses and pseudo-aligned inputs as training strategies. We adopt the interpretable and evaluation-friendly approach [35] of warping the ground truth to the input using OF, allowing accurate pixel-wise supervision and metric computation.
Hyperspectral datasets. Several datasets support HSR research. Early datasets like CAVE [37] and Harvard [38] provided controlled hyperspectral measurements. Later datasets such as ICVL [17], KAUST [39], and ARAD_1K [23] focused on enabling data-driven HSR methods. These datasets contain either radiance or reflectance data, but most lack real RGB images and instead simulate RGB via CMFs, which fails to capture the characteristics of camera pipelines. Moreover, they often assume perfect alignment, which does not hold in practical settings. Recent datasets like BeyondRGB [5] and MobileSpec [4] address some of these issues by including real RGB captures. BeyondRGB includes color charts and lightsource spectrum estimation, while MobileSpec offers aligned RGB-HSI pairs. However, they still face trade-offs between alignment, diversity, and color reference availability.We build on these efforts by introducing Doomer dataset with real RGB images, misaligned hyperspectral data, and in-scene color references thus enabling realistic reconstruction under natural capture conditions.
@@
Doomer
143 Radiance Gray ball
@@
Table 1: Comparison of existing hyperspectral datasets. Our proposed Doomer dataset uniquely offers real multi-view RGB images with spectral filters, misalignment, and in-scene gray reference under diverse conditions.
@@ 3.1 System Overview
Our goal is to improve HSR by increasing the number of spectrally distinct measurements captured simultaneously, using only consumer-grade hardware. Rather than relying on custom sensor arrays or coded optics, we build upon multi-camera smartphones, which are already equipped with multiple rear-facing cameras featuring different lenses and spectral sensitivity functions (SSFs).
To amplify spectral diversity, we augment the auxiliary cameras with external filters. As illustrated in Fig. 1, this converts each RGB camera into a spectrally modulated sensor. The resulting device captures nine distinct spectral channels: three from each RGB camera, modified by its filter without requiring scanning or internal hardware changes.
Formally, let $\mathbf{S}_{i}\in\mathbb{R}^{3\times n}$ be the SSF of camera $i$, and $\mathbf{f}_{i}\in[0,1]^{n}$ the spectral transmittance of the filter applied to that camera. The effective per-camera response becomes $\mathbf{S}_{i}\circ\mathbf{f}_{i}$, and the overall system response is:
$\mathbf{S}_{\mathbf{F}}=\begin{bmatrix}\mathbf{S}_{1}\circ\mathbf{f}_{1}\\ \mathbf{S}_{2}\circ\mathbf{f}_{2}\\ \ldots\\ \mathbf{S}_{k}\circ\mathbf{f}_{k}\end{bmatrix}\in\mathbb{R}^{3k\times n},$
where $\mathbf{F}=[\mathbf{f}_{1}^{\top};\dots;\mathbf{f}_{k}^{\top}]$ denotes the filter configuration. In our prototype, $k=3$, with one unfiltered camera ($\mathbf{f}_{1}=\mathbf{1}$) and two filtered cameras. This design simplifies dataset collection (Sec. 4) while still significantly expanding the observed spectral basis.
@@ 3.2 Filter Selection via Spectral Uncertainty Minimization
Given a library of $N=65$ candidate filters available in our lab, we exhaustively evaluate all $65\times 64$ ordered filter pairs for the two auxiliary cameras. The optimal pair should minimize spectral ambiguity — i.e., the uncertainty in the latent spectrum $\mathbf{r}\in\mathbb{R}^{n}$ given a measurement $\mathbf{c}=\mathbf{S}_{\mathbf{F}}\mathbf{r}+\mathbf{n}$, where $\mathbf{n}$ is sensor noise.
Following [40], we use the expected conditional variance of the spectrum as our selection criterion:
$v(\mathbf{F})=\mathbb{E}_{\mathbf{c}}\left[\mathrm{tr}\,\mathbb{V}_{\mathbf{r}}(\mathbf{r}\mid\mathbf{c})\right].$
This criterion reflects the average spectral uncertainty remaining after observing $\mathbf{c}$. Lower $v(\mathbf{F})$ implies more informative measurements and improved reconstructability.
Figure 2: Filters selected for Tele and Wide cameras.
To compute this metric, we sample spectra $\mathbf{r}$ from a uniform distribution over pixels in a publicly available hyperspectral dataset [39] (licensed under CC BY-NC 3.0). We use precise SSFs for our smartphone cameras and spectrophotometer-measured filter transmittances (see Sec. A.1 for details). The final filter pair (Fig. 2) is selected as the one that minimizes $v(\mathbf{F})$.
@@ 4 Doomer Dataset
Overview. To support training and evaluation for the MI-HSR task introduced in Section 3, we collect a new dataset, which we call Doomer [^1] The name Doomer is inspired by the subcultural aesthetic: most scenes were collected under gloomy or overcast weather conditions, in contrast to the brightly lit existing datasets.. Existing hyperspectral datasets are not suited for our setup: none provide spatially misaligned multi-view RGB observations. Furthermore, no existing dataset aligns with our specific hardware configuration (a triple-camera smartphone with custom spectral filters), making new data collection a necessity.
Doomer contains 143 real-world scenes captured using a Huawei Mate 40 Pro smartphone equipped with Main, Tele, and Wide cameras, along with a Specim IQ hyperspectral camera for ground truth. In each scene, we record three smartphone RAW images, two with custom spectral filters and one unfiltered as well as a 111-band (400 – 730 nm) hyperspectral image. Example captures are shown in Figure 3.
Scenes include a mix of indoor and outdoor environments under varying illumination conditions (e.g., halogen and LED lighting indoors; overcast and sunny weather outdoors), and span a range of objects including food, printed material, architectural surfaces, and color calibration charts. Each scene includes a gray ball reference for future work on illumination estimation. While the gray ball is visible in Wide and Main cameras, it falls outside the Tele field of view and is cropped out during preprocessing. Non-preprocessed versions will be included in the public release.
Figure 3: Sample scenes from the Doomer dataset. Smartphone images are rendered to sRGB using device-specific color matrices; hyperspectral images are rendered using CIE RGB CMF.
@@
Capture setup. The data acquisition rig consists of a Specim IQ hyperspectral camera, a Huawei Mate 40 Pro smartphone mounted in a 3D-printed case with slots for spectral filters and a gray reference sphere (VFX ball).
The entire system is mounted on a tripod, with the phone positioned to rotate along a vertical axis for alternating captures (Figure 4). This design allows the smartphone and hyperspectral camera to image scenes from nearly identical viewpoints, minimizing parallax and occlusion. The gray ball is connected to a rigid rod that allows to regulate the position of the ball in the scene.
@@
Preprocessing pipeline. Each four-image scene group (three RGB + one HSI) undergoes standardized preprocessing (see Supp.). We also normalize field of view and resolution across sensors. Specifically, we estimate pairwise homographies between each RGB image and the Tele view using SIFT keypoints and RANSAC. All images are then warped to the Tele frame and cropped accordingly (Fig. 5). When automatic registration fails, manual alignment is used. Despite this correction, residual geometric misalignment remains due to parallax and non-planar scene structure — motivating our use of alignment-aware HSR models. Finally, all images are downsampled to match the hyperspectral resolution, originally $194\times 259$ cropped to $192\times 256$ for convenience, and spectral grid is resampled to conventional 400–730 nm, $n=34$, to optimize computational costs.

### v4

@@ Abstract
Hyperspectral reconstruction (HSR) from RGB images is a highly promising direction for accurate color reproduction and material color measurement. While most existing approaches rely on a single RGB image — thereby limiting reconstruction accuracy — the majority of modern smartphones are equipped with two or more cameras. In this work, we propose a novel multi-image-to-hyperspectral reconstruction (MI-HSR) framework that leverages a triple-camera smartphone system, where two lenses are equipped with carefully selected spectral filters. Our easy-to-implement configuration, based on theoretical and empirical analysis, allows to obtain more complete and diverse spectral data than traditional single-chamber setups. To support this new paradigm, we introduce Doomer, the first dataset for MI-HSR, comprising aligned images from three smartphone cameras and a hyperspectral reference camera across diverse scenes. We further introduce a lightweight alignment module for MI-HSR that effectively fuses multi-view inputs while mitigating parallax- and occlusion-induced artifacts. Proposed module demonstrate consistent quality improvements for modern HSR methods. In a nutshell, our setup allows 30% more accurate estimations of spectra compared to an ordinary RGB camera, while the proposed alignment module boosts the reconstruction quality of SotA methods by an additional 5%. Our findings suggest that spectral filtering of multiple views with commodity hardware unlocks more accurate and practical hyperspectral imaging.
@@ 1 Introduction
Hyperspectral imaging provides dense spectral measurements at each spatial pixel, forming a 3D cube ${\bm{\mathsfit{Y}}}\in\mathbb{R}^{h\times w\times n}$, where $n\gg 3$. This enables fine-grained analysis of material properties in applications ranging from remote sensing [41], to medical diagnostics [37], to historical preservation [21], to ISP improvement [48, 18], to food quality assessment [31, 2]. However, acquiring such high-dimensional data typically requires hardware that is expensive, bulky, and often reliant on time-consuming scanning thus fundamentally limiting the usability of hyperspectral imaging in dynamic or consumer settings.
@@
An increasingly studied alternative is hyperspectral reconstruction (HSR): recovering $\bm{\mathsfit{Y}}$} from an RGB observation ${\bm{\mathsfit{X}}}\in\mathbb{R}^{h\times w\times 3}$ captured via a sensor with a spectral sensitivity matrix ${\bm{C}}\in\mathbb{R}^{3\times n}$ under noise realization $\bm{\mathsfit{N}}$}:
${\bm{\mathsfit{X}}}={\bm{\mathsfit{Y}}}{\bm{C}}+{\bm{\mathsfit{N}}}.$
@@
This inverse problem is highly ill-posed and despite advances in deep learning-based HSR [5, 47], reconstructing $\bm{\mathsfit{Y}}$} from a single RGB view remains fundamentally limited by the low spectral observability.
@@
Several efforts have aimed to improve this observability. Learned or optimized multispectral filter arrays (MSFAs) [40], end-to-end spectral sensitivity learning [26], and joint sensor-network co-design [16] have all been proposed. However, these methods assume control over sensor hardware and manufacturing, thus making them impractical for scalable or consumer-level deployment.
@@
One underexplored but promising path is to leverage the multi-camera systems found in modern smartphones. These devices already include multiple rear cameras with different lenses and spectral sensitivities. In principle, such a setup can be treated as a low-cost, multi-spectral capture system thus capable of observing a scene through multiple, distinct sensitivity matrices. However, prior works on leveraging multiple cameras [29, 27] do not tackle unavoidable image misalignment.
@@
Can we turn a multi-camera smartphone into a compact, spectrally diverse imaging system, without altering its internal hardware? Our key insight is that by modulating the spectral response of the auxiliary cameras using carefully chosen external filters, we can create an imaging system that allows richer spectral representations. Unlike synthetic MSFAs or custom hardware, our approach requires no internal modification and enables real-time, parallel acquisition. Critically, this setup is easy to deploy, manufacturable at scale, and compatible with existing mobile infrastructure.
@@
We select filters using spectral information loss minimization with respect to a prior hyperspectral distribution. The resulting setup produces spatially misaligned but spectrally rich multi-view data thus posing a new fusion problem that we address via alignment-aware learning. An overview of our physical configuration is shown in Fig. 1, which illustrates how external filters are applied to the auxiliary lenses of a standard smartphone to create a spectrally diverse input set. The combination of low-cost physical augmentation and learning-based reconstruction represents a practical path toward deployable hyperspectral imaging in unconstrained environments. We propose a complete, low-cost pipeline for multi-image-to-hyperspectral reconstruction (MI-HSR) using a filter-modified smartphone. Specifically, the contributions are three-fold:
  • A novel, low-cost spectral acquisition system that converts a commodity triple-camera smartphone into a 9-channel imaging device. This is achieved by augmenting the auxiliary cameras with two external spectral filters, which are theoretically selected using a spectral uncertainty minimization criterion to maximize captured information.
  • A novel framework for multi-image hyperspectral reconstruction (MI-HSR), featuring our novel alignment-aware block. It fuses spatially misaligned and spectrally diverse inputs by leveraging optical-flow-guided deformable convolutions. Such utilization of auxiliary inputs allows +2.44 dB PSNR improvement over single-image HSR.
  • The Doomer dataset, the first real-world benchmark for the MI-HSR task. The dataset contains 155 real-world scenes, each providing spatially misaligned RAW images from the filtered smartphone setup and a corresponding ground-truth hyperspectral scan.
@@ 2 Related Work
Doomer 155 Radiance Gray ball
@@
Table 1: Overview of existing datasets for HSR from RGB. Our proposed Doomer dataset uniquely offers real multi-view RGB images with spectral filters, misalignment, and in-scene gray reference under diverse conditions.
Low-cost multispectral imaging
Numerous approaches have aimed to capture multispectral or hyperspectral information without expensive hardware. Early work by Helling et al. [19] employed a grayscale camera with a rotating filter wheel, while Valero et al. [36] used an RGB camera with three interchangeable filters. Oh et al. [27] captured scenes with three different cameras, leveraging the variation in their spectral sensitivities. More recently, Sharma et al. [29] demonstrated that consumer mobile devices with both RGB and NIR sensors can achieve extended spectral capture (400–1000 nm). Although these systems reduce costs, they either involve long capture times or assume no misalignment between successive captures.
Hyperspectral reconstruction from RGB
Traditional HSR methods model spectra using sparse coding [3], dictionary learning [1], or manifold embeddings [20], based on the low-dimensional structure of hyperspectral data and the rarity of metamers [14]. These approaches are computationally efficient but often lack the capacity to incorporate global context from the input image, making them less robust to complex natural scenes. Recent methods based on deep learning have achieved significant advances, particularly those developed through the NTIRE spectral reconstruction challenges [4]. Early deep models used CNNs [17, 30], while newer transformer-based approaches like MST++ [5] and SPECAT [43] introduced attention along spectral or spatial dimensions, and diffusion based approaches like SDM [42] or LDERT [24] also adapted for HSR. However, nearly all research relies on synthetic RGB inputs rendered from hyperspectal images (HSIs) using color matching functions (CMFs), assuming perfect alignment and sometimes access to camera parameters. These assumptions do not hold in practice, limiting model generalizability and non-sustainable to hard metameric cases [15, 42]. We address these limitations training and evaluating on real-world data with acquisition artifacts and misalignments.
Hyperspectral datasets
Early datasets like CAVE [44] and Harvard [6] provided controlled hyperspectral measurements. Later datasets such as ICVL [3], KAUST [25], and Arad 1K [4] focused on enabling data-driven HSR methods. These datasets contain either radiance or reflectance data, but most lack real RGB images and instead simulate RGB via given sensor spectral sensitivities, which fails to capture the characteristics of camera pipelines. Moreover, they often assume perfect alignment, which does not hold in practical settings. Recent datasets like BeyondRGB [18] and MobileSpec [48] address some of these issues by including real RGB captures. BeyondRGB includes color charts and light source spectrum estimation, while MobileSpec offers aligned RGB-HSI pairs. However, they still face trade-offs between alignment, diversity, and color reference availability. We build on these efforts by introducing Doomer dataset with real misaligned RGB images, hyperspectral data, and in-scene color references thus enabling realistic reconstruction under natural capture conditions.
Handling misalignment
Misalignment between inputs is a well-known challenge in video and reference-based super-resolution tasks. Optical flow (OF) [22, 7], deformable convolutions [34, 39], attention mechanisms [38] have been proposed to mitigate this. In our context, misalignment also arises due to ground truth HSIs being not spatially aligned with the RGB input. Zhang et al. [46] addressed this by warping ground truth toward the input using OF, enabling pixel-level evaluation. Elezabi et al. [11] proposed contextual losses and pseudo-aligned inputs as training strategies. We adopt the interpretable and evaluation-friendly approach [46] of warping the ground truth to the input using OF, allowing accurate pixel-wise supervision and metric computation.
@@ 3 Proposed Imaging System
Figure 2: Sample scenes from the Doomer dataset. Smartphone images are rendered to sRGB using color matrices from RAW metadata; hyperspectral images are rendered using CIE RGB color matching functions.
@@ 3.1 System Overview
To amplify spectral diversity, we cover Tele- and Wide cameras with external filters. As illustrated in Fig. 1, this converts each RGB camera into a spectrally modulated sensor. The resulting device captures nine distinct spectral channels without requiring scanning or hardware changes.
Formally, let ${\bm{C}}_{i}\in\mathbb{R}^{3\times n}$ be the spectral sensitivity functions (SSF) of camera $i$, and $\bm{f}_{i}\in[0,1]^{n}$ the spectral transmittance of the filter applied to that camera. Thus, per-camera response becomes ${\bm{C}}_{i}\odot{\bm{f}}_{i}$, and the overall response is:
${\bm{C}}_{{\bm{F}}}=\begin{bmatrix}({\bm{C}}_{0}\odot{\bm{f}}_{0})^{\top};\ldots;({\bm{C}}_{k-1}\odot{\bm{f}}_{k-1})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{3k\times n},$
where ${\bm{F}}=[{\bm{f}}_{0}^{\top};\dots;{\bm{f}}_{k-1}^{\top}]$ is the filter configuration and $\odot$ is element-wise multiplication. In our prototype, $k=3$, with one unfiltered camera (${\bm{f}}_{0}=(1,\dots,1)^{\top}$) and two filtered cameras. This design simplifies dataset collection (Sec. 4) while still significantly enriching spectral representations.
@@ 3.2 Filter Selection via Spectral Uncertainty Minimization
Given a library of $65$ candidate filters from standard set, we exhaustively evaluate all $65\times 64$ ordered filter pairs for the two auxiliary cameras. The optimal pair should minimize spectral ambiguity — i.e., the uncertainty in the latent spectrum ${\mathbf{y}}\in\mathbb{R}^{n}$ given a measurement ${\mathbf{x}}={\bm{C}}_{{\bm{F}}}{\mathbf{y}}+{\mathbf{n}}$, where $\mathbf{n}$} is sensor noise.
Following [28], we use the expected conditional variance of the spectrum as a selection criterion:
$v({\bm{F}})=\mathbb{E}_{{\mathbf{x}}}\left[\mathrm{tr}\,\mathrm{Var}_{{\mathbf{y}}}({\mathbf{y}}\mid{\mathbf{x}})\right].$
This criterion reflects the average spectral uncertainty remaining after observing $\mathbf{x}$}. Lower $v({\bm{F}})$ implies more informative measurements and improved reconstructability.
To compute this metric, we sample spectra $\mathbf{y}$ from a uniform distribution over pixels in a publicly available hyperspectral dataset [25]. We use precise SSFs for our smartphone cameras and spectrophotometer-measured filter transmittances (see Sec. A.1 for details). The final filter pair (Fig. 9 in Supp.) is selected as the one that minimizes $v({\bm{F}})$.
To verify effectiveness of the proposed filter selection strategy, we trained the MI-HSR method in simulated settings with 12 different filter sets. See Sec. A.2 for details.
@@ 4 Doomer Dataset
Overview
To support training and evaluation for the MI-HSR task introduced in Sec. 3, we collect a new dataset, which we call Doomer [^1] The name Doomer is inspired by the subcultural aesthetic: most scenes were collected under gloomy or overcast weather conditions, in contrast to the brightly lit existing datasets.. Existing hyperspectral datasets are not suited for our setup: none provide spatially misaligned multi-view real RGB observations. Furthermore, no existing dataset aligns with our specific hardware configuration (a triple-camera smartphone with custom spectral filters), making new data collection a necessity.
Doomer contains 155 real-world scenes captured using a Huawei Mate 40 Pro smartphone equipped with Main, Tele, and Wide cameras, along with a Specim IQ hyperspectral camera for ground truth. In each scene, we record three smartphone RAW images, two with custom spectral filters and one unfiltered as well as a 111-band (400 – 730 nm) hyperspectral image. Example captures are shown in Fig. 2.
To provide representability, scenes include a mix of indoor and outdoor environments under varying illumination conditions (e.g., halogen and LED lighting indoors; overcast and sunny weather outdoors), and span a range of objects including food, printed material, architectural surfaces, and color calibration charts. Each scene includes a gray ball reference for future work on illumination estimation. While the gray ball is visible in Wide and Main cameras, it falls outside the Tele field of view and is cropped out during preprocessing. Non-preprocessed RAW data versions will be included in the public release.
@@
Capture setup
The data acquisition rig consists of a Specim IQ hyperspectral camera, a Huawei Mate 40 Pro smartphone mounted in a 3D-printed case with slots for spectral filters and a gray reference sphere (VFX ball).
The entire system is mounted on a tripod, with the phone positioned to rotate along a vertical axis for alternating captures (Fig. 3). This design allows the smartphone and hyperspectral camera to image scenes from nearly identical viewpoints, minimizing parallax and occlusion. The gray ball is connected to a rigid rod that allows to regulate the position of the ball in the scene.
@@
Data preprocessing
Each four-image scene group (three RGB + one HSI) undergoes standardized preprocessing (see Sec. A.3). We also normalize field of view and resolution across sensors. Specifically, we estimate pairwise homographies between each RGB image and the Tele view using SIFT keypoints and RANSAC. All images are then warped to the Tele frame and cropped accordingly. When automatic registration fails, manual alignment is used. Despite this correction, residual geometric misalignment remains due to parallax and non-planar scene structure — motivating our use of alignment-aware HSR models. Finally, all images are downsampled to match the hyperspectral resolution, originally $194\times 259$ cropped to $192\times 256$ for convenience, and spectral grid is resampled to conventional 400–730 nm, $n=34$, to optimize computational costs.